OpenStudy (anonymous):

Solve Linear Equation and simplify: -2/5(z-4)=2/5(z-3)

OpenStudy (midhun.madhu1987):

Cross Multiply first..

OpenStudy (anonymous):

Excuse me if im wrong: (z-4)=(z-3) ??

OpenStudy (midhun.madhu1987):

you did a good try.. but there is a slight mistake.. u missed a Negative sign.. hope u understood..

OpenStudy (anonymous):

-(z-4)=(z-3) ?

OpenStudy (midhun.madhu1987):

-(z-3)=(z-4)

OpenStudy (midhun.madhu1987):

now solve for z.

OpenStudy (anonymous):

z=7/2 ?

OpenStudy (anonymous):

7/2 is correct..